\begin{matrix}a + b + c & uv & x − y & 27 \\ a + b & u + v & z & 134 \end{matrix}

(\begin{matrix}x + y + z & xy \\ 2x + xy & xz \end{matrix}) (\begin{matrix}223 & 43 \\ 34 & 15 \end{matrix})

[\begin{matrix}a + b + c & uv & : & 27 \\ a + b & u + v & : & 134 \end{matrix}]

|\begin{matrix}5 & 3 & 4 & 2 \\ 1 & 3 & 6 & 4 \end{matrix}|

∥\begin{matrix}5 & 3 & 4 & 2 \\ 1 & 3 & 6 & 4 \end{matrix}∥